OpenStudy (anonymous):

This question is complex to write so I hope everyone understands. Find f^-1(x). f(x) = Cube Root 3\/x-9

OpenStudy (anonymous):

solve \[x=\sqrt[3]{y-9}\] for \(y\)

OpenStudy (anonymous):

cube both sides to get \[x^3=y-9\] then add 9 to both sides and you are done

OpenStudy (anonymous):

y = x^3+9

OpenStudy (anonymous):

Is that:\[y=\sqrt[3]{x-9}\]Since the inverse has the range of the original function as its domain and the domain of the original function as its range, we switch the y and x:\[\rightarrow x = \sqrt[3]{y-9}\] Now you have to solve for y and get the equation in the form, y = .... Can you do that? @MixMasterTX

OpenStudy (anonymous):

And yes that's correct. Good job. @MixMasterTX

OpenStudy (anonymous):

So when you cube both sides it cancels out the cube root on the right side?

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

Taking both sides to the power/exponent of 3 makes the cube root go away. Then you can solve for y quite easily.

OpenStudy (anonymous):

ok...so the f^-1(x) really plays no part in the equation?

OpenStudy (mertsj):

Yes it does. Didn't you notice right at the beginning that Satellite formed the inverse by switching the x and the y?

OpenStudy (anonymous):

Saying f(x) = something is the same as saying y = something. It just means that f is a function of x or y is changing with respect to x. So that way we can replace f(x) with y when we are finding inverses so it's easier and makes more sense when you graph it on the xy plane. @MixMasterTX

OpenStudy (anonymous):

ok...so by switching the x and the y it allows you to cube both sides thus eliminating the cube on the right side, adding 9 to both sides and solving. y = x^3+9......Right?

OpenStudy (anonymous):

Thanks to @genius12 @satellite73 @Mertsj

OpenStudy (mertsj):

No. By switching the x and the y, you are forming the inverse.